How many 4-letter code words are possible using the first 8 letters of the alphabet when letters are allowed to repeat?
I know the answer is 4096, but what is the answer when you cannot repeat letters? and how did you get given answer?

Answer by sudhanshu_kmr(1152)


no. of possible code words when repetition allowed = 8 * 8 * 8 * 8 = 4096

no. of possible code, when repetition is not allowed = 8P4 = 1680
as repetition is not allowed,
no. of ways to arrange 4 letters out of 8 = 8P4
another ways,
no. of ways to choose first letter for code = 8
no. of ways to choose second letter for code = 7 (except the first one )
similarly, no. of ways to choose third and fourth letter are 6 and 5.
total no. of ways = 8*7*6*5 = 1680